OpenStudy (anonymous):

I need help with this question: A car completes a journey at an average speed of 40 km/h. At what speed must it travel on the return journey if the avg. speed for the complete journey (out and back) is 60 km/h? (I'm confused with the question, so explain it step by step please)

OpenStudy (ahsome):

Here is what it is saying: \[\text{A car travels from point A to B, and back to B}\]\[\text{The speed he drives at from Point A to B is 40km/h}\]\[\text{The average speed he drives from Point A to B, and back to A is 60km/h}\] Do you understand that, @zGrimReaper123 ?

OpenStudy (ahsome):

\[\text{They want you to find the speed from Point B to Point A}\]

OpenStudy (ahsome):

Do you, @zGrimReaper123?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

But can you explain it bit by bit? I'm not really great at maths

OpenStudy (ahsome):

Its fine. I just want to know if you understand the above completely?

OpenStudy (ahsome):

If you don't understand ANYTHING, just say so.

OpenStudy (anonymous):

I don't understand anything

OpenStudy (ahsome):

Ok, Imagine there is a car. He is going from one place to another. He is going at a speed of 40km/h

OpenStudy (ahsome):

Lets say the first place is Point A, the second place is Point B

OpenStudy (anonymous):

I get that part

OpenStudy (anonymous):

I'm just unsure of how I'm supposed to calculate the whole thing

OpenStudy (ahsome):

Ok. Have you heard of this equation: \[Average=\frac{Total 1 + Total 2 + Total 2+ Total 4...}{\text{Number of Totals}}\]

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Does it mean average speed is equal to total distance divided by total time?

OpenStudy (ahsome):

Yes, The issue there is that we are not given the time OR distance :(

OpenStudy (ahsome):

So we must try another way

OpenStudy (anonymous):

Right

OpenStudy (ahsome):

Good. Now you know these informatin: \[\text{The First Total(Speed)}\]\[\text{How many totals there are (2)}\] Now, we don't know the speed of the car that is travelling back, so let that be \(x\)

OpenStudy (anonymous):

Ok...

OpenStudy (ahsome):

Why are you confused?

OpenStudy (anonymous):

I'm not, I'm just agreeing

OpenStudy (anonymous):

To your point

OpenStudy (ahsome):

Oh, phew ;)

OpenStudy (anonymous):

Yeah hehe

OpenStudy (ahsome):

Average Equation:\[Average=\frac{Total1+Total2+Total2+Total4...}{Number of Totals}\] Sub in the Average Speed, The First Speed, The Second Speed, and the number of totals: \[60km/h=\frac{40km/h+x}{2}\] Remove the the \(km/h\) \[60=\frac{40+x}{2}\]

OpenStudy (ahsome):

Do you understand that?

OpenStudy (anonymous):

What does sub stand for?

OpenStudy (ahsome):

Substitue in. Pretty much, put information inside.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

I mean, I already know the answer is 120, I just want to know the method, and so far you're really helpful

OpenStudy (ahsome):

Now, lets find what \(x\) is, which we said was the speed of the return trip \[60km=\frac{40km+x}{2}\] Now to remove the 2 on the right hand side, we need to do the opposite of what the 2 is doing, In this equation, it is dividing. Therefore, we need to multiply both sides by 2. \[60*2=\frac{40+x}{2}*2\] \[120=40+x\] Now, move the 40 to the left hand side. Since in this example it is adding 40, we need to take away 40 on both sides. \[120-40=40+x-40\] \[80=x\]\[x=80\] Therefore, the speed when he came back was \(80km/h\)

OpenStudy (ahsome):

Does that make sense, @zGrimReaper123?

OpenStudy (anonymous):

Yes, but are you sure the answer is 80 km/h?

OpenStudy (ahsome):

According to my calculations yeah, why?

OpenStudy (ahsome):

@hartnn, if you can clarify on the question please?

OpenStudy (anonymous):

The question I'm talking about was taken from a book I'm using, and the answer for it is written at the back, which is 120km/h, not 80km/h

OpenStudy (anonymous):

Finally my pc decided to send it, it was hanging

OpenStudy (ahsome):

Hmm. From my calculations, it cannot be. However, I might be missing something...

OpenStudy (anonymous):

Yeah, probably

OpenStudy (ahsome):

It can't be though. Since the 60 is the average speed, and the equation: \[Average=\frac{Total1+Total2...}{\text{Number of Totals}}\] If we were to sub in the values 40, 2 and 120(Your answer) \[60=\frac{40+120}{2}\] \[60=\frac{160}{2}\] \[60=80\] Thats impossible. Either I missed some crucial information, or the book is wrong.

OpenStudy (cj49):

40+80=120/2 wud give u an avg speed of 60 but 40+120=160/2 wud give u and avg speed of 80 so i guess@Ahsome is right the ans is 80 and not 120

hartnn (hartnn):

average speed = total distance / total time

OpenStudy (ahsome):

We don't know distance OR time. So what we do?

hartnn (hartnn):

assume some variables

OpenStudy (anonymous):

@Ahsome Perhaps you're right. I wrote down the entire question, so probably your calculations are correct after all

OpenStudy (anonymous):

Or maybe not

OpenStudy (ahsome):

Thanks @zGrimReaper123. If I figure out the difference, I will tell you. I am guessing they forgot to take away 40, which gave 120 as an answer.

OpenStudy (anonymous):

I guess so

hartnn (hartnn):

A -> B Speed = 40km/hr Distance = D time = D/40 hrs. A->B->A Speed = 60 km/hr Distance = 2D time = 2D/60 hrs

hartnn (hartnn):

B->A Speed = x Distance = D time = D/x find x

hartnn (hartnn):

60 = 2D/ (D/40 + D/x)

hartnn (hartnn):

60 = 2/ (1/40+1/x) find x

OpenStudy (ahsome):

Is this it? \[60=\frac{2}{\frac{1}{40}+\frac{1}{x}}\]

hartnn (hartnn):

you will indeed get x=120 which is correct

hartnn (hartnn):

the catch here was to apply the formula avg. speed = total distance / total time total distance = D+D total time = time from A->B + time from B->A = D/40 +D/x

OpenStudy (ahsome):

Cross Multiply? \[60*(\frac{1}{40}+\frac{1}{x})=2*1\]

hartnn (hartnn):

yep

OpenStudy (ahsome):

\[\frac{60}{240}+\frac{60}{60x}=2\]

hartnn (hartnn):

simpler way to do it will be 1/x = 1/30 -1/40

OpenStudy (ahsome):

Get everything to have the same base \[\frac{60}{240}*\frac{x}{x}+\frac{60}{60x}*\frac{4}{4}=\frac{2}{1}*\frac{240}{240}\]

OpenStudy (ahsome):

\[\frac{60x}{240x}+\frac{240}{240x}=\frac{420}{240x}\] Cut the \(240x\)

OpenStudy (ahsome):

\(60x+240=480\) \(60x=240\) \(x=4\)

OpenStudy (ahsome):

Now what?

hartnn (hartnn):

wot ? O.o what have you done ?

OpenStudy (ahsome):

I must have made some HUGE mistake. Anyway ;)

OpenStudy (ahsome):

I see where I made a mistake, I also times the denominator ;)

hartnn (hartnn):

60 (1/40 +1/x) = 2 1/40 +1/x = 2/60 = 1/30 1/x = 1/30 - 1/40 = (40-30)/1200 = 10/1200 = 1/120 x=120

OpenStudy (ahsome):

hopefully @zGrimReaper123 understood all this ;)

hartnn (hartnn):

important thing is to get this equation avg speed = total distance / total time 60 = 2D/ (D/40 + D/x) then its pretty algebraic.

OpenStudy (ahsome):

Yes, @hartnn, yes it is ;)

OpenStudy (ahsome):

Do you need us to reexplain that, @zGrimReaper123?

OpenStudy (anonymous):

@Ahsome Maybe, just let me read the explanation again. I was away while you guys were at it

OpenStudy (anonymous):

@Ahsome @hartnn I really appreciate your help :D

OpenStudy (ahsome):

No problem. Thank @hartnn, I went on some tangent ;)

OpenStudy (anonymous):

no i have an easy way to solve this

OpenStudy (ahsome):

:D

OpenStudy (anonymous):

u are so good ahsome

OpenStudy (anonymous):

how are u

OpenStudy (anonymous):

@TheGreatStupidPerson Please tell me the easier way

OpenStudy (anonymous):

you can ask your father if he is good at studies

OpenStudy (anonymous):

ahsome is your grandmother alive

OpenStudy (anonymous):

@TheGreatStupidPerson Are you fluttering serious?

OpenStudy (anonymous):

what saying me not understandings

OpenStudy (anonymous):

no teej me ok

OpenStudy (anonymous):

i will complaint your teacher if u teej me

OpenStudy (anonymous):

now leave your studies and let have a break

OpenStudy (anonymous):

@TheGreatStupidPerson Can you piss off somewhere else?

OpenStudy (anonymous):

for enjoyment i will tell u a story of a princess ok

OpenStudy (anonymous):

@Ahsome

OpenStudy (anonymous):

@Ahsome Are we able to chat privately?

OpenStudy (anonymous):

you can ask your teacher if she knows the correct answer

OpenStudy (anonymous):

@zGrimReaper123 why is your name say that u are a raper and u rape women